OpenStudy (anonymous):

Complete the standard form of the equation 3y/2=8-x ___x+___y=16

OpenStudy (texaschic101):

3y/2 = 8 - x --- multiply everything by 2 3y = 16 - 2x --- add 2x to both sides 2x + 3y = 16

OpenStudy (anonymous):

Thnx

OpenStudy (texaschic101):

sure thing :)